OpenStudy (anonymous):

Find the axis of symmetry of the graph of the function. f(x) = 2x2 - 16x + 27

OpenStudy (anonymous):

would that occur at the minimum point of the quadratic?

OpenStudy (dls):

vertex maybe

OpenStudy (dls):

just use the formula..if i'm right

OpenStudy (john_es):

Of course, for both.

OpenStudy (dls):

The vertex of a quadratic equation is the minimum or maximum point of the equation.

OpenStudy (anonymous):

reckon we should find the deriviative of f(x) and get stationary point

OpenStudy (jim766):

x = -b/2a

OpenStudy (jim766):

a = 2 b = -16

OpenStudy (dls):

yeah^^

OpenStudy (dls):

y=c-b^2/4a

OpenStudy (jim766):

the axis is syemetry, the vertical line that divides a parabola (the graph of a quadratic) is found by using the formula x = -b/2a in this case x = 16/2 or x = 8

OpenStudy (anonymous):

shouldnt it be x=4

OpenStudy (anonymous):

is the axis of symmetry

OpenStudy (jim766):

yes, my bad, I forgot to mult denom by 2 x = 16/4 or x = 4 thanks, good catch!

OpenStudy (anonymous):

np.